Sodium carbonate and silver nitrate react to form silver carbonate and sodium nitrate.
Write a balanced chemical equation for the reaction.
If 3.00 grams of sodium carbonate reacts with 5.00 grams of silver nitrate, what is the limiting reactant?
What is the excess reactant?
How many grams of silver carbonate are produced?
How many grams of sodium nitrate are produced?
How many grams of the excess reactant are left over?
1
Expert's answer
2020-05-01T14:42:51-0400

Na2CO3(aq)+2AgNO3(aq) -->Ag2CO3(s)+2NaNO3(aq);

1.n(Na2CO3)=3/106=0.0283; n(AgNO3)=5/(2*170)=0.0147.

 Silver nitrate is the limiting reactant,

Sodium carbonate is the excess reactant

2.m(Ag2CO3)=0.0147*Mr(Ag2CO3)=0.0147*276=4 g

3.m(NaNO3)=0.0147*2*85=2.5 g

4.m(Na2CO3)=(0.0283-0.0147)*106=1.44 g